OpenStudy (anonymous):

determine if the improper integral of 4/sin(x)(sqrt(1-x^2) from 0 to 1converges and if so at what value?

OpenStudy (loser66):

\[\int_0^1\frac{4}{sinx \sqrt{1-x^2}}~~~or~~~\int_0^1 \frac{4}{sinx}\sqrt{1-x^2}dx\] which one?

OpenStudy (anonymous):

the first one

OpenStudy (loser66):

I am sorry, I really don't know,

OpenStudy (anonymous):

thats ok neither do i

OpenStudy (anonymous):

this is very complicated stuff

OpenStudy (loser66):

@ash2326

OpenStudy (psymon):

\[\frac{ 4 }{ \sin(x)\sqrt{1-x ^{2}} }\]? What loser tried to post shows as an error for me, so just want to confirm if this is it

OpenStudy (anonymous):

yes also you take the integral of that from 0 to 1

OpenStudy (inkyvoyd):

maybe integration by parts and trig sub?

OpenStudy (anonymous):

how so?

OpenStudy (inkyvoyd):

well it looks like your u and v' will be 1/sin x and 1/sqrt(1-x^2)

OpenStudy (inkyvoyd):

you knnow how to find the integral of both expressions, but you really want v to be what I chose. try it out and lmk how it goes; I have to do homework myself unfortunately

OpenStudy (psymon):

Yeah, that does work out actually. I guess I was just double checking myself.

OpenStudy (inkyvoyd):

btw that's my excuse for not doing the problem since I don't feel like it lol

OpenStudy (anonymous):

ok i'll try that thanks

OpenStudy (psymon):

Mhm. You'll have to do integration by parts twice. It comes out pretty clean, though.

OpenStudy (inkyvoyd):

protip if you're allowed to use a table for the trig sub, it's a pain in the arse

OpenStudy (inkyvoyd):

@Psymon want a even simpler integration problem that's really hard to evaluate?

OpenStudy (psymon):

Lol, why would I want to do more unnecessary work? xD This one just turned out to be okay.

OpenStudy (inkyvoyd):

does that integral even converge btw?>

OpenStudy (psymon):

Nope.

OpenStudy (inkyvoyd):

LOL

OpenStudy (psymon):

But just saying that isnt enough, he still has to be able to work his way through it to determine that, lol.

OpenStudy (inkyvoyd):

yeah guy42 use limits

OpenStudy (anonymous):

it doesn't converge? are you sure?

OpenStudy (psymon):

Positive.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

nope it does converge

OpenStudy (inkyvoyd):

gg

OpenStudy (psymon):

Really? O.o

OpenStudy (anonymous):

the not converge answer choice was wrong

OpenStudy (inkyvoyd):

do you get another try?

OpenStudy (anonymous):

3 more tries

OpenStudy (psymon):

I dont see how, haha. \[\int\limits_{0}^{1}\frac{ 4 }{ \sin(x)\sqrt{1-x ^{2}} } \] If thats the integral then Im trying to see how that actually converges x_x

OpenStudy (inkyvoyd):

psymon go to wolfram alpha, and someone gimme a medal or inky be sad

OpenStudy (psymon):

wolfram goes all dead trying to do it

OpenStudy (inkyvoyd):

kay lemme try mathematica

OpenStudy (psymon):

Oh wow.....I think I got an answer x_x This is one funky integral, questionable accuracy like crazy.

OpenStudy (anonymous):

i got the answer (pi^2)/2

OpenStudy (psymon):

And thats still not the answer I got, haha. Like seriously, wtf?

OpenStudy (inkyvoyd):

ok, did you guys try out what I suggested with the by parts and all?

OpenStudy (psymon):

Yeah, I did.

OpenStudy (inkyvoyd):

does it work?

OpenStudy (psymon):

I just gotta follow through with it. I guess I need to see if I can trust my pencil and paper more than wolfram

OpenStudy (anonymous):

ok guys im gonna go thanks for your help

OpenStudy (psymon):

e

OpenStudy (inkyvoyd):

if you take v'=csc x you get Log[Sin[x/2]] Log[Tan[x/4]] + Log[1 - I Tan[x/4]] Log[Tan[x/4]] + Log[1 + I Tan[x/4]] Log[Tan[x/4]] - 1/2 Log[Tan[x/4]]^2 + Log[(-(1/2) - I/2) (-1 + Tan[x/4])] Log[-I + Tan[x/4]] + Log[(-(1/2) + I/2) (-1 + Tan[x/4])] Log[I + Tan[x/4]] + Log[-I + Tan[x/4]] Log[(1/2 - I/2) (1 + Tan[x/4])] + Log[I + Tan[x/4]] Log[(1/2 + I/2) (1 + Tan[x/4])] - Log[Tan[x/4]] Log[1 + Tan[x/4]] - 1/2 Log[Sin[x/2]] Log[1 - Tan[x/4]^2] + 1/2 Log[Tan[x/4]] Log[1 - Tan[x/4]^2] - 1/2 Log[-I + Tan[x/4]] Log[1 - Tan[x/4]^2] - 1/2 Log[I + Tan[x/4]] Log[1 - Tan[x/4]^2] - 1/2 Log[Sin[x/2]] Log[-1 + Tan[x/4]^2] + 1/2 Log[Tan[x/4]] Log[-1 + Tan[x/4]^2] - 1/2 Log[-I + Tan[x/4]] Log[-1 + Tan[x/4]^2] - 1/2 Log[I + Tan[x/4]] Log[-1 + Tan[x/4]^2] + PolyLog[2, 1 - Tan[x/4]] + PolyLog[2, 1/2 ((1 + I) - (1 - I) Tan[x/4])] - PolyLog[2, -Tan[x/4]] + PolyLog[2, -I Tan[x/4]] + PolyLog[2, I Tan[x/4]] + PolyLog[2, (-(1/2) - I/2) (I + Tan[x/4])] + PolyLog[2, 1/2 ((1 + I) + (1 - I) Tan[x/4])] + PolyLog[2, 1/2 ((1 - I) + (1 + I) Tan[x/4])] - (Log[ Cos[x/2]] (Log[2] Log[1 - Tan[x/4]] - 1/2 Log[1 - Tan[x/4]]^2 + Log[Cos[x/4]^2] Log[Tan[x/4]] + Log[1 - I Tan[x/4]] Log[Tan[x/4]] + Log[1 + I Tan[x/4]] Log[Tan[x/4]] + Log[(-(1/2) - I/2) (-1 + Tan[x/4])] Log[-I + Tan[x/4]] + Log[(-(1/2) + I/2) (-1 + Tan[x/4])] Log[I + Tan[x/4]] + Log[-I + Tan[x/4]] Log[(1/2 - I/2) (1 + Tan[x/4])] + Log[I + Tan[x/4]] Log[(1/2 + I/2) (1 + Tan[x/4])] + Log[2] Log[1 + Tan[x/4]] - 2 Log[1 - Tan[x/4]] Log[1 + Tan[x/4]] - 1/2 Log[1 + Tan[x/4]]^2 - 1/2 Log[Cos[x/4]^2] Log[1 - Tan[x/4]^2] - 1/2 Log[-I + Tan[x/4]] Log[1 - Tan[x/4]^2] - 1/2 Log[I + Tan[x/4]] Log[1 - Tan[x/4]^2] - 1/2 Log[Cos[x/4]^2] Log[-1 + Tan[x/4]^2] - 1/2 Log[-I + Tan[x/4]] Log[-1 + Tan[x/4]^2] - 1/2 Log[I + Tan[x/4]] Log[-1 + Tan[x/4]^2] - PolyLog[2, 1/2 (1 - Tan[x/4])] + PolyLog[2, 1/2 ((1 + I) - (1 - I) Tan[x/4])] - PolyLog[2, -Tan[x/4]] + PolyLog[2, -I Tan[x/4]] + PolyLog[2, I Tan[x/4]] - PolyLog[2, Tan[x/4]] + PolyLog[2, (-(1/2) - I/2) (I + Tan[x/4])] - PolyLog[2, 1/2 (1 + Tan[x/4])] + PolyLog[2, 1/2 ((1 + I) + (1 - I) Tan[x/4])] + PolyLog[2, 1/2 ((1 - I) + (1 + I) Tan[x/4])]))/(Log[ Cos[x/4]^2] + Log[1 - Tan[x/4]] + Log[1 + Tan[x/4]])

OpenStudy (psymon):

Whoops. Okay, it does converge, nvm x_x

OpenStudy (inkyvoyd):

what am I not seeing sigh

OpenStudy (inkyvoyd):

maybe u-sub?

OpenStudy (psymon):

Yeah, it converges, I just needed to do the actual work, lol.

OpenStudy (inkyvoyd):

wait take x=cos theta

OpenStudy (inkyvoyd):

wait no that doesn't work

OpenStudy (inkyvoyd):

maybe we could use power series to reduce it into something more manageble?

OpenStudy (psymon):

You wanna see it?

OpenStudy (psymon):

I have an answer.

OpenStudy (inkyvoyd):

hell yes

OpenStudy (psymon):

Now of course catch me if you see something I did wrong, haha.

OpenStudy (psymon):

Bleh, nvm, saw an error in what I was doing -_- lame.

OpenStudy (inkyvoyd):

this is one of those problems that I don't think are solveable

OpenStudy (psymon):

Im too lazy to try trig sub, lol.

OpenStudy (inkyvoyd):

no dude, I feel like it's not expressible in terms of elementary functions

OpenStudy (psymon):

Probably not. Ive gotten 50 different answers fromdifferent methods. And then none of them matched up with what he came up with. Meh, ohwell.

OpenStudy (inkyvoyd):

dude, this problem is on magnitude of difficulty of the basel problem

OpenStudy (inkyvoyd):

*I think

OpenStudy (inkyvoyd):

how do you know this integral converges btw